# Developer Perspectives on Licensing and Copyright Issues Arising from Generative AI for Coding

### v1

@@ 5.2.1. Sentiments regarding the inclusion of one’s own code in training data
Of the 554 respondents familiar with GenAI tools, the largest proportion (40.6%), as seen in Figure 11, indicated that they would be neither pleased nor displeased if the code they developed was included in training data without their permission. However, open-ended responses indicate that the reality is context-dependent. If the work was open-sourced, respondents generally indicated that they would be pleased or even flattered to have their code included in training data as a way in which they could contribute to the greater good. $R_{144}$ expressed this sentiment: “If it’s just a class, or a few elements from my code, I would actually be pleased and proud that my code is having [an] impact on other peoples’ work and on the advancement of science in general.” $R_{67}$ elaborated, saying, “If the code has been used [in training data] with my implicit (e.g., through BSD-licensed code) or explicit (e.g., through my company’s or my permission) then I would be happy to have contributed. Otherwise [my] answer becomes ‘displeased.’ ” (Some respondents, however, expressed concerns about training models on open-source code given the quality of the training data. $R_{196}$ stated, “[M]ost of the code I write is bad or middling, and the lack of intentional design I’ve seen in the code of open-source projects […] really worries me […] [in that] we can just automatically generate unclear and potentially insecure code.”)
@@
Some developers were also concerned that the unauthorized use of work as training data for GenAI would disincentivize human creativity, consistent with the traditional justification for copyright law in the United States. If users increasingly seek solutions from LLMs rather than directly from the original project, developers may be reluctant to continue to contribute to the open-source community, causing the well of open-source training data to eventually run dry. As $R_{148}$ put it, “[This] is more of a case of how to sustain OS[S], than [it is] a case of code being copied.” Indeed, $R_{245}$ said, if developers are not incentivized to create code, “we will end up with a gradual decline in the quality of all intellectual property as current AI systems are not creative enough to produce IP that is significantly different from the training material.” $R_{514}$ offered a similar view: “The real issue to me is that those engines are [parasiting off] websites like [S]tack [O]verflow [and] actually lowering the amount of money they make. If these websites stop […] produc[ing] good answers that [AI tools] can feed [off of], they [the AI tools] all will die.” $R_{604}$ even speculated that “[c]ode generation tools might encourage creative professionals to hoard knowledge, defeating the original purpose of copyright: enabling protected[,] limited-time monetization followed by release to the public domain.”
@@ 5.2.2. Conditions around use of one’s code in training data
Some respondents argued that the inclusion of proprietary code in training data warranted monetary compensation. Relatedly, many believed that the developers of open-source software had waived their right to any compensation. For some respondents, the answer depended on whether the owner of the AI model profited from the use of the data. $R_{159}$, for example, stated, “[I]f my code is free, and the AI model is paid, I feel wronged and deserve compensation.” $R_{613}$ offered a similar view: “[T]he developers whose code these big corporations use to train their models must receive some amount of compensation in terms of free credits, [GPU] time[,] etc. on their platform.”
@@
$R_{433}$ took a broader view: “If monetary compensation were required to train these models, it would put smaller AI development companies out of business and restrict the currently thriving sector to only the Tech Giants. No one else would have the [means] to not only acquire the data but also to do the training.”
@@
Finding 13: Some respondents suggested that training data contributors should receive monetary compensation or other benefits, such as free access to the trained model or free credits for the platform. However, forcing AI companies to provide monetary compensation to all contributors of the training data, while seemingly fair, would place an additional hurdle potentially blocking smaller AI companies from entering the market.
@@
For many respondents, it was also important for licensing terms to be respected, including attribution for their contributions or (for restrictive licenses) reuse under the same terms. Modes of attribution, as respondents suggested, could be as fine-grained as a citation provided along with the generated code or as coarse as a manifest that included links to all the training data. In the view of respondents, embedding attribution into the AI model would serve two functions: (1) respecting the rights of developers and (2) helping users and organizations ensure that they are complying with the licenses of code they use by providing provenance of the output. As to the latter, $R_{456}$ imagined a situation where an organization using GenAI unintentionally reproduced GPL-licensed code verbatim, introducing license compliance issues into the project. At least two respondents felt that responsibility for avoiding such problems rested with the models. $R_{378}$ said, “The onus is on the LLM creators to make sure [the output] includes the source” because “[t]he users are not at fault[…].” $R_{665}$ noted, similarly, “Developers should not bear the burden for […] copyright […] concerns. Tools should not be producing code which could get a developer in trouble.”
@@
Some respondents would favor the creation of new OSS licenses that provide specific conditions for inclusion in training data. Such licenses would enable creators to control how and when their works are used for GenAI. As $R_{587}$ put it: “I think that open source developers that hold strong opinions about the use of their code with AI models need to update their licenses to enforce those desires.”
@@
5.2.3. Fair use and replicated code
While some developers, as described above, believed that the use of code as training data requires a license, others considered such use to be a fair use, even if it was not always clear whether respondents were referring to this concept in its legal sense or in a more general sense. For example, $R_{173}$ said, “I believe that AI training utilizing any published content is a form of fair use and should be encouraged rather than litigated. The usage is entirely transformative rather than derivative.” $R_{695}$ said, “I think training on copyrighted materials is legal, but their outputs can violate [copyright law] surprisingly easily (even worse when considering trademarks). I’ve had them reproduce IDENTICAL text to copyrighted sources, which is clearly infringement. AI companies want to not have liability for such infringement, putting it on the users, but that’s stupid.” For some, the extent of the copying was important, perhaps reflecting U.S. copyright law’s concept of de minimis copying. $R_{144}$ put it this way: “I feel like if significant pieces of code are copied, only then does it really become an issue for me: I would not be comfortable if someone took multiple files from my code and used them in a proprietary project.” $R_{143}$ said, “If a generative AI happens to have copied large swathes of code, yes [the developers of that code deserve attribution].”
Some respondents who believed that the use of content as training data was fair use compared the practice to the process through which humans produce new works based on the knowledge they have accumulated over time. These respondents therefore believed that it should be assumed that code made publicly available is free to use for training, and the burden should be on developers to indicate otherwise. For example, $R_{26}$ said, “Creators need to make efforts to protect data that should not be used in AI training.” This was echoed by $R_{96}$: “If your product is a unique algorithm, then it is your responsibility to keep it closed” and by $R_{429}$: “If an individual or institution wishes to have their data removed from a training dataset, they should have the right to request its removal.” (These respondents did not suggest how developers might accomplish such removal.)
In follow-up interviews, some respondents elaborated that they attempt to exercise such control by limiting the types of projects for which they use GenAI tools, thereby restricting the tools’ access to their code. $RI_{118}$ specified that when choosing tools to use, “I’m not going to use a tool at work that’s going to expose my work to risk […], whereas if it’s just a personal project, I’m a lot less concerned. I want to make sure I’m not going to violate anything major or there’s nothing truly weird like they’re going to own my code.”
Finding 15: Some developers believed that the use of code as training data should be considered fair use by default and that the onus should be on developers to protect code that they do not wish to be included in training data.
@@ 5.3. AI-generated output that resembles existing code
Developers also had a range of views on whether it is wrongful for generated code to closely resemble existing code. Some respondents compared what GenAI models do when producing code to what humans have always done (manually) in creating derivative work. $R_{103}$ stated: “Being an open-source software developer, I’m accustomed to seeing my codes copied or ‘cloned’ by other developers without any notice, and I don’t feel anything special from it.” $R_{177}$ asserted, “All developers copy code.” $R_{265}$ noted, “What AI is doing now was being done by developers [throughout] the years through forums, [S]tack [O]verflow, [G]ithub, etc.” $R_{130}$ remarked, “Most developers had been using Stack Overflow and OS[S] projects for years for similar purposes before AI became widely available.”
Finding 16: A group of developers did not see how GenAI technology was doing anything significantly different from traditional means of creating code, which reuses existing code created by others. The only distinction is the scale and ease by which it can be done.
For other respondents, their view on the wrongfulness of generated code that resembled code they had developed depended on the nature and amount of the code being copied, an intuition that tracks U.S. legal doctrine (17 U.S.C. §102, b). $R_{687}$ said that, in their view, wrongfulness depends on “if [the] result is a simple hello world vs. word for word of an entire codebase [ ] In my current workflow, expected outputs are only snippets, so I am indifferent about legality and attribution.” This was an attitude shared by other respondents. $R_{183}$ stated that “small pieces of code are rarely valuable in themselves. For any given function there’s only so many ways to write it. It’s the entire thing that’s valuable.” $R_{246}$ provided a concrete example, saying, “There are only so many ways you can create a fetch request in Java[S]cript,” and $R_{90}$ put it this way: “[T]o say setting up code to talk to a server or render UI is unique is an exaggeration. Even your design, is what[,] a three column layout? That is not someone’s to own. It’s all been done a million times already[,] and we are all reusing founders technology going back to punch cards.” $R_{573}$ summarized the concept: “Development is typically a process of piecing together various snippets that are customized to a particular use case; there is no expectation of ‘ownership,’ at least in the case of partial code blocks. While the reproduction of an entire application/code base for use in generating revenue is problematic, the sharing of common code is how the development community has progressed [until] now.” And $R_{445}$ offered a useful analogy: “[A]dding a license to a particular piece of code does not make sense. You can add licenses to libraries and exact forms of algorithms. Adding a license to a piece of code is like adding a copyright to individual strokes and colors in a painting.”
@@ 5.4. On the ownership of AI-generated code
Our discussion to this point has focused on developers’ views on potential liability for copying another’s code, either in training data or in the generated output. But another important legal question is who would own any copyright to code created by GenAI models. In this subsection, we do not attempt to provide a legal analysis of this question; rather, our goal is to understand what developers think the answer to the question is.
Participants were presented with a list of potential options as to who owns the copyright to code generated through the use of AI and were invited to select all that they believed applied, with the opportunity to provide other options and to further explain their answers. Even though courts have yet to resolve this question, only 57 of 554 respondents familiar with GenAI tools (10.3%) indicated that they were not sure who owned the output or had no opinion on the matter. The rest of the respondents expressed an opinion and reported high confidence in their answers: 416 (75.1%) stated they were confident or even very confident.
@@ 5.4.1. No one—works are in the public domain
The most popular choice, selected by 242 participants (43.7%), was that AI-generated code belongs to no one. Some, including $R_{7}$, reasoned this way because they viewed the AI itself as the true author of the code: “The code belongs to the author—the AI model. It has no rights, so its work must be in the public domain.” $R_{199}$ stated, similarly, “The generated code can’t have any copyright as [it] wasn’t made by a person.” $R_{516}$ offered a summary: “Code generated through the use of AI is not copyrightable, because it fails the creative step required by law of being a product of the intellect. Instead it is produced by a non-sentient machine, lacks a creative step, an expression of the soul. It is therefore in the public domain.”
@@
Several respondents invoked the inherently collaborative process of generating code through AI as a reason not to assign copyright ownership to only one individual or entity. $R_{74}$ gave an analogy: “It’d be like saying who gets the copyright of what a child creates. [T]heir teachers? [T]heir parents? [T]heir friends? [T]hey all influence it in an unpredictable way. At least to my knowledge[,] we can’t measure it.” Putting it in more concrete terms, $R_{105}$ stated that the output of GenAI models is “a product of complex interactions between AI algorithms, training data, and user inputs, making traditional copyright assignment impractical. This perspective aligns with the open-source ethos, favoring communal access and innovation over individual ownership in the rapidly evolving field of AI and software development.” $R_{518}$ said that the process of training and using LLMs “mirrors how all software development builds upon the foundation of previous programs and code […], reflecting the collaborative nature of programming itself.”
This sentiment regarding the nature of collaborative work also was echoed in responses, particularly from open-source developers, that seemed to suggest that even if the developer could claim copyright in AI output, they should refrain from asserting their rights in favor of the greater good. $R_{233}$ asserted that the focus should be on maximizing utility: “I’ve been contributing to open-source code for over a decade, consistently opting for the least restrictive licenses. I firmly believe in the value of sharing knowledge rather than imposing limitations. With the emergence of AI as a lasting tool, my stance is to focus on maximizing its utility for people rather than getting bogged down in minor details.” $R_{546}$ said, similarly, “[W]hen I publish open-source code I desire that this code is re-used and helps as many people as possible. If AI models can scale this further and democratize the use of open-source code even more, I see this as highly desirable.” $R_{695}$ concluded, “[I]f we’re going to destroy a bunch of people’s livelihoods, to produce something I find rather uncompelling, having it contribute to society as a whole via public domain would be great.”
Finding 18: The largest group of developers (242, 43.68%) thought that the output of GenAI belongs in the public domain. Some believe this because it is consistent with the collaborative nature of open-source software development, while others believe that, because AI models are not human, the output of those models cannot be subject to copyright law.
@@ 5.4.2. The prompting developer or their employer
The next most popular suggestion was that model output belonged to the developer that prompted the model or their employer, as selected by 201 respondents (36.3%).
@@
Many participants who reached this conclusion believed that the use of code-generation tools is similar to existing processes that can result in copyrightable code. Respondents with this perspective, like $R_{173}$, characterized GenAI as “a tool like any other, used to augment the skills and direction of its user,” where the assumption would be that the user owned the rights to whatever was created using the tool. $R_{22}$ provided an analogy: “[I]f a carpenter chooses to use a powerful automation tool like [a] CNC machine to make something, the final product is his (or his employer’s), not the maker of the tool or [whoever] inspired the carpenter.” $R_{236}$ summarized the sentiment as follows: “When one uses a tool to create something, you should own the result of your work.” Ultimately, $R_{118}$ indicated, “If I can copyright code I learned by essentially copying what I learned from Stack Overflow, I should be able to copyright what I get from […] Copilot. […] If I can’t, then the analysis should be similar. The focus on AI as the issue seems misguided. The only difference is scale.”
@@
For many developers, the prompt was an important starting point for the ownership question. $R_{129}$ stated plainly, “Without the prompt, there is no output.” $R_{187}$ explained, “[C]ode generation requires prompt work, and the copyright on the code that is generated should be the result of that prompt work.” For other developers, the claim to ownership relied on the assumption that code resulting from GenAI is not ready to use immediately upon generation but must be further refined by the developer. As explained by $R_{107}$, “AI code generation typically doesn’t result in a perfectly working code. It’s a sketch, or draft variant that will be modified and fine[-]tuned […]. AI doesn’t do development on its own[;] it helps developers do their job.” $R_{63}$ concluded that if the models were capable of “creating a large amount of code without [requiring] editing,” copyright issues might be more complex, but $R_{63}$ was “not convinced that AI code generation is good enough for that.” $R_{39}$ stated similarly, “The development process of using AI development tools usually involves AI generating an incorrect outline of one snippet of code in a larger system, that I then amend, correct and test. The additional work I put into it seems like it should constitute a new creative work that is copyrightable.”
@@
Finding 19: The second largest group of developers (201, 36.3%) believed that the user of a GenAI tool should own the copyright to code generated with that tool because it is the user’s contributions—creating the prompt and, typically, editing and further refining the result—that create the code. In other words, GenAI was seen as simply a new tool for creating content.
@@
One challenge raised by some respondents, however, is that even if those supplying the prompts are deemed to be the owners of copyright in the resulting code, the nature of code creation with GenAI is such that similar code might result from two independently created prompts. Such activities could result, under U.S. law, in two separate copyrights, one for each generated work. $R_{98}$ found this legal conclusion to be counterintuitive: “[T]he developers who prompted can’t also be the copyright owners – it’s fairly simple to get very similar or identical outputs from [a] LLM by two different developers.” $R_{276}$ expressed a similar concern, stating, “I believe that the prompt was the thing that caused the generation of the given code. And it could be generated again with a similar prompt,” as did $R_{507}$, who wrote, “If I prompt the model to write code that I later sell, can I copyright it? Yes of course. […] Can someone else prompt it the same way and maybe get the exact same code? Yes. Does that make sense? No. Which makes me think that AI copyright makes no sense.”
@@
Ultimately, some respondents noted, control over the resulting work incentivizes creative activity. $R_{300}$ asserted, “AI will not be a useful product to sell if the authors of the original training data or the model creators retain copyright on generated stuff, or if the output is uncopyrightable.” $R_{114}$ mentioned that they “specifically use [c]ode [g]eneration tools that include in their ToS specific statements that copyright belongs to us (the customer),” and $R_{288}$ stated, “If the license does not end up with me at the end of the circus, I will use a different tool made by a different company trained off different data.”

### v3

@@ 5.2.1. Sentiments regarding the inclusion of one’s own code in training data
Of the 554 respondents familiar with GenAI tools, the largest proportion (41%), as seen in Figure 8, indicated that they would be neither pleased nor displeased if the code they developed was included in training data without their permission. However, open-ended responses indicate that the reality is context-dependent. If the work was open-sourced, respondents generally indicated that they would be pleased or even flattered to have their code included in training data as a way in which they could contribute to the greater good. $R_{144}$ expressed this sentiment: “If it’s just a class, or a few elements from my code, I would actually be pleased and proud that my code is having [an] impact on other peoples’ work and on the advancement of science in general.” $R_{67}$ elaborated, saying, “If the code has been used [in training data] with my implicit (e.g., through BSD-licensed code) or explicit (e.g., through my company’s or my permission) then I would be happy to have contributed. Otherwise [my] answer becomes ‘displeased.’ ” (Some respondents, however, expressed concerns about training models on open-source code given the quality of the training data. $R_{196}$ stated, “[M]ost of the code I write is bad or middling, and the lack of intentional design I’ve seen in the code of open-source projects […] really worries me […] [in that] we can just automatically generate unclear and potentially insecure code.”)
@@
Consistent with the traditional justification for copyright law in the United States, $R_{148}$ reasoned that unauthorized use of work as training data for GenAI could disincentivise human creativity, leading to reluctance from developers to contribute to open source: “[This] is more of a case of how to sustain OS[S], than [it is] a case of code being copied.” Indeed, $R_{245}$ said, if developers are not incentivized to create code, “we will end up with a gradual decline in the quality of all intellectual property as current AI systems are not creative enough to produce IP that is significantly different from the training material.” $R_{514}$ offered a similar view: “The real issue to me is that those engines are [parasiting off] websites like [S]tack [O]verflow [and] actually lowering the amount of money they make. If these websites stop […] produc[ing] good answers that [AI tools] can feed [off of], they [the AI tools] all will die.” $R_{604}$ even speculated that “[c]ode generation tools might encourage creative professionals to hoard knowledge, defeating the original purpose of copyright: enabling protected[,] limited-time monetization followed by release to the public domain.”
@@ 5.2.2. Conditions around use of one’s code in training data
Eleven respondents utilized the open-ended questions in the survey to provide their thoughts on whether developers should receive monetary compensation for their code being included in training data. Four of these argued that training on open-source software did not necessitate such compensation or, indeed, that the developers of open-source software had waived their right to any compensation. Six identified some conditions under which they viewed it to be appropriate. $R_{159}$, for example, stated, “[I]f my code is free, and the AI model is paid, I feel wronged and deserve compensation.” Lastly, $R_{613}$ believed that developers should always receive some form of compensation: “[T]he developers whose code these big corporations use to train their models must receive some amount of compensation in terms of free credits, [GPU] time[,], etc. on their platform.”
@@
Separate from this group, $R_{433}$ took a broader view: “If monetary compensation were required to train these models, it would put smaller AI development companies out of business and restrict the currently thriving sector to only the Tech Giants. No one else would have the [means] to not only acquire the data but also to do the training.”
@@
Finding 13: Some respondents suggested that training data contributors should receive monetary compensation or other benefits, such as free access to the trained model or free credits for the platform. However, requiring AI companies to provide monetary compensation to all contributors to the training data could make it more difficult for smaller AI companies to enter the market.
@@
Eight respondents stated that it was also important to respect licensing terms, including attribution for their contributions or (for restrictive licenses) reuse under the same terms. As respondents suggested, modes of attribution could be as fine-grained as a citation provided along with the generated code or as coarse as a manifest that included links to all the training data. In the view of respondents, embedding attribution into the AI model would serve two functions: (1) respecting the rights of developers and (2) helping users and organizations ensure that they comply with the licenses of code they use by providing provenance of the output. As to the latter, $R_{456}$ imagined a situation where an organization using GenAI unintentionally reproduced GPL-licensed code verbatim, introducing license compliance issues into the project. At least two respondents felt responsible for avoiding such problems rested with the models. $R_{665}$ noted, “Developers should not bear the burden for […] copyright […] concerns. Tools should not be producing code which could get a developer in trouble.”
@@
Another avenue would be for developers to create or choose new OSS licenses that provide specific conditions for inclusion in training data. Such licenses would enable creators to control how and when their works are used for GenAI. As $R_{587}$ put it: “I think that open source developers that hold strong opinions about the use of their code with AI models need to update their licenses to enforce those desires.”
@@
5.2.3. Fair use and training data
While some developers, as described above, believed that the use of code as training data requires a license, others expressed in open-ended responses that they considered such use to be a fair use, even if it was not always clear whether respondents were referring to this concept in its legal sense or a more general sense. For example, $R_{173}$ said, “I believe that AI training utilizing any published content is a form of fair use and should be encouraged rather than litigated. The usage is entirely transformative rather than derivative.”
Some respondents who believed that using content as training data was fair use compared the practice to the process through which humans produce new works based on the knowledge they have accumulated over time. These respondents believed that it should be assumed that publicly available code is free to use for training, and the burden should be on developers to indicate otherwise. For example, $R_{26}$ said, “Creators need to make efforts to protect data that should not be used in AI training.”
In follow-up interviews, one respondent elaborated that they attempt to exercise such control by limiting the types of projects for which they use GenAI tools, thereby restricting the tools’ access to their code. $RI_{118}$ specified that when choosing tools to use, “I’m not going to use a tool at work that’s going to expose my work to risk […], whereas if it’s just a personal project, I’m a lot less concerned. I want to make sure I’m not going to violate anything major or there’s nothing truly weird like they’re going to own my code.”
Finding 15: Some developers believed that using code as training data should be considered fair use by default and that the onus should be on developers to protect code that they do not wish to be included in training data.
@@ 5.3. AI-generated output that resembles existing code
Developers also had a range of views on whether it is wrongful for generated code to closely resemble existing code. Some respondents compared what GenAI models do when producing code to what humans have always done (manually) in creating derivative work. $R_{103}$ stated: “Being an open-source software developer, I’m accustomed to seeing my codes copied or ‘cloned’ by other developers without any notice, and I don’t feel anything special from it.” $R_{265}$ noted, “What AI is doing now was being done by developers [throughout] the years through forums, [S]tack [O]verflow, [G]ithub, etc.”
Finding 16: A group of developers did not see how GenAI technology was doing anything significantly different from traditional means of creating code, which reuses existing code created by others.
For other respondents, their view on the wrongfulness of generated code that resembled code they had developed depended on the nature of the code being copied, an intuition that tracks U.S. legal doctrine (17 U.S.C. §102, b). $R_{183}$ stated that “small pieces of code are rarely valuable in themselves. For any given function there’s only so many ways to write it. It’s the entire thing that’s valuable.” $R_{246}$ provided a concrete example, saying, “There are only so many ways you can create a fetch request in Java[S]cript.” $R_{573}$ summarized the concept: “Development is typically a process of piecing together various snippets that are customized to a particular use case; there is no expectation of ‘ownership,’ at least in the case of partial code blocks. While the reproduction of an entire application/code base for use in generating revenue is problematic, the sharing of common code is how the development community has progressed [until] now.” And $R_{445}$ offered a useful analogy: “[A]dding a license to a particular piece of code does not make sense. You can add licenses to libraries and exact forms of algorithms. Adding a license to a piece of code is like adding a copyright to individual strokes and colors in a painting.”
For some, the extent of the copying was important, perhaps reflecting U.S. copyright law’s concept of de minimis copying, which does not constitute infringement. $R_{144}$ put it this way: “I feel like if significant pieces of code are copied, only then does it really become an issue for me: I would not be comfortable if someone took multiple files from my code and used them in a proprietary project.” $R_{143}$ said, “If a generative AI happens to have copied large swathes of code, yes [the developers of that code deserve attribution].”
@@ 5.4. On the ownership of AI-generated code
Our discussion to this point has focused on developers’ views on potential liability for copying another’s code, either in training data or in the generated output. Another important legal question is who would own any copyright to code created by GenAI models. In this subsection, we do not attempt to provide a legal analysis of this question; rather, our goal is to understand what developers think the answer to the question should be.
Participants were presented with a list of potential options as to who owns the copyright to code generated through the use of AI and were invited to select all that they believed applied, with the opportunity to provide other options and to further explain their answers. Even though courts have yet to resolve this question, only 57 of 554 respondents familiar with GenAI tools (10%) indicated that they were not sure who owned the output or had no opinion on the matter. The rest of the respondents expressed an opinion and reported high confidence in their answers: 416 (75%) stated they were confident or even very confident.
@@ 5.4.1. No one—works are in the public domain
The most popular choice, selected by 242 participants (44%), was that AI-generated code belongs to no one. $R_{516}$ offered a summary: “Code generated through the use of AI is not copyrightable, because it fails the creative step required by law of being a product of the intellect. Instead it is produced by a non-sentient machine, lacks a creative step, an expression of the soul. It is therefore in the public domain.”
@@
Several respondents invoked the inherently collaborative process of generating code through AI as a reason not to assign copyright ownership to only one individual or entity. $R_{74}$ gave an analogy: “It’d be like saying who gets the copyright of what a child creates. [T]heir teachers? [T]heir parents? [T]heir friends? [T]hey all influence it in an unpredictable way. At least to my knowledge[,] we can’t measure it.” Putting it in more concrete terms, $R_{105}$ stated that the output of GenAI models is “a product of complex interactions between AI algorithms, training data, and user inputs, making traditional copyright assignment impractical. This perspective aligns with the open-source ethos, favoring communal access and innovation over individual ownership in the rapidly evolving field of AI and software development.”
This sentiment regarding the nature of collaborative work was also echoed in responses, particularly from open-source developers, that seemed to suggest that even if the developer could claim copyright in AI output, they should refrain from asserting their rights in favor of the greater good. $R_{233}$ asserted that the focus should be maximizing utility: “I’ve been contributing to open-source code for over a decade, consistently opting for the least restrictive licenses. I firmly believe in the value of sharing knowledge rather than imposing limitations. With the emergence of AI as a lasting tool, my stance is to focus on maximizing its utility for people rather than getting bogged down in minor details.”
Finding 18: The largest group of developers (242, 44%) thought that the output of GenAI belongs in the public domain. Some believe this because it is consistent with the collaborative nature of open-source software development, while others believe that because AI models are not human, the output of those models cannot be subject to copyright law.
@@ 5.4.2. The prompting developer or their employer
The next most popular suggestion was that model output belonged to the developer that prompted the model or their employer, as selected by 201 respondents (36%).
@@
Many participants who reached this conclusion believed that using code-generation tools is similar to existing processes that can result in copyrightable code. Respondents with this perspective, like $R_{173}$, characterized GenAI as “a tool like any other, used to augment the skills and direction of its user,” where the assumption would be that the user owned the rights to whatever was created using the tool. $R_{22}$ provided an analogy: “[I]f a carpenter chooses to use a powerful automation tool like [a] CNC machine to make something, the final product is his (or his employer’s), not the maker of the tool or [whoever] inspired the carpenter.” Ultimately, $R_{118}$ indicated, “If I can copyright code I learned by essentially copying what I learned from Stack Overflow, I should be able to copyright what I get from […] Copilot. […] If I can’t, then the analysis should be similar. The focus on AI as the issue seems misguided. The only difference is scale.”
@@
Six developers explicitly stated that the prompt was an important starting point for the ownership question. $R_{187}$ explained, “[C]ode generation requires prompt work, and the copyright on the code that is generated should be the result of that prompt work.” For other developers, the claim to ownership relied on the assumption that code resulting from GenAI is not ready to use immediately upon generation but must be further refined by the developer. As explained by $R_{107}$, “AI code generation typically doesn’t result in a perfectly working code. It’s a sketch, or draft variant that will be modified and fine[-]tuned […]. AI doesn’t do development on its own[;] it helps developers do their job.” $R_{63}$ concluded that if the models were capable of “creating a large amount of code without [requiring] editing,” copyright issues might be more complex, but $R_{63}$ was “not convinced that AI code generation is good enough for that.” $R_{39}$ stated similarly, “The development process of using AI development tools usually involves AI generating an incorrect outline of one snippet of code in a larger system that I then amend, correct and test. The additional work I put into it seems like it should constitute a new creative work that is copyrightable.”
@@
Finding 19: The second largest group of developers (201, 36%) believed that the user of a GenAI tool should own the copyright to code generated with that tool because it is the user’s contributions—creating the prompt and, typically, editing and further refining the result—that create the code. In other words, GenAI was seen as simply a new tool for creating content.
@@
One challenge raised by respondents, however, is that even if those supplying the prompts are deemed to be the owners of copyright in the resulting code, the nature of code creation with GenAI is such that similar code might result from two independently created prompts. Such activities could result, under U.S. law, in two separate copyrights, one for each generated work. $R_{98}$ found this legal conclusion to be counterintuitive: “[T]he developers who prompted can’t also be the copyright owners—it’s fairly simple to get very similar or identical outputs from [a] LLM by two different developers.” $R_{507}$ wrote, “If I prompt the model to write code that I later sell, can I copyright it? Yes of course. […] Can someone else prompt it the same way and maybe get the exact same code? Yes. Does that make sense? No. Which makes me think that AI copyright makes no sense.”
@@
Ultimately, some respondents noted that control over the resulting work incentivizes creative activity. $R_{300}$ asserted, “AI will not be a useful product to sell if the authors of the original training data or the model creators retain copyright on generated stuff, or if the output is uncopyrightable.” $R_{114}$ mentioned that they “specifically use [c]ode [g]eneration tools that include in their ToS specific statements that copyright belongs to us (the customer),” and $R_{288}$ stated, “If the license does not end up with me at the end of the circus, I will use a different tool made by a different company trained off different data.”